Example for seminar.sty

Policarpa Salabarrieta July 21, 1991

Information overload = “Too much” information

You have 134 unread messages:
Do you want to read them now?

1. People cannot process all the information they receive.

2. People should receive less information.

University of Guaduas, March 13, 1998 1-a

Information overload = “Too much” information

You have 134 unread messages:
Do you want to read them now?

1. People cannot process all the information they receive.

2. People should receive less information.

University of Guaduas, March 13, 1998 1-b

Information overload = “Too much” information

You have 134 unread messages:
Do you want to read them now?

1. People cannot process all the information they receive.

2. People should receive less information.

• There is information overload in a network if there is some mechanism that, compared to the status quo, makes the senders and/or receivers better off by restricting the flow of information.’

• There is information overload in a network if there is some mechanism that, compared to the status quo, makes the senders and/or receivers better off by restricting the flow of information.

• There is information overload in a network if there is some mechanism that, compared to the status quo, makes the senders and/or receivers better off by restricting the flow of information.’

• There is information overload in a network if there is some mechanism that, compared to the status quo, makes the senders and/or receivers better off by restricting the flow of information.

Questions

• When could there be overload in networks?

• What mechanims make the receivers and senders better off?

• How does the welfare of the senders and receivers depend on the cost of communication?

Questions

• When could there be overload in networks?

• What mechanims make the receivers and senders better off?

• How does the welfare of the senders and receivers depend on the cost of communication?

University of Guaduas, March 13, 1998 3-a-red

Questions

• When could there be overload in networks?

• What mechanims make the receivers and senders better off?

• How does the welfare of the senders and receivers depend on the cost of communication?

University of Guaduas, March 13, 1998 3-a-blue

Questions

• When could there be overload in networks?

• What mechanims make the receivers and senders better off?

• How does the welfare of the senders and receivers depend on the cost of communication?

Being more informed

is always better, but it’s not the same as

receiving more information Being more informed

is always better, but it’s not the same as

receiving more information

A tax τ on communication is said to support X̃(c) if X̃(c) is an equilibrium for Γ(c + τ).

Proposition 6. Assume X̃(c) is not an equilibrium for Γ(c).

1. If supp(γ) = [0, 1]^{n}, there is no tax that supports X̃(c).

2. If supp(γ) = S^{n−1}, there is a tax that supports X̃(c) if and only if m = 1, p_{j} > c ∀j, and

(a) n = 2; or

(b) n = 3 and p_{i}^{−1} + p_{j}^{−1} ≥ p_{k}^{−1} for all distinct i, j, k; or

(c) n = 4 and p_{1} = p_{2} = p_{3} = p_{4}.

Architecture

[figure: 1; 2; 3; 4; 5; 6; 7; 8]